Hi! Pretty please with a cherry on the top would you help me solve this problem! Thank you so much. Appreciate it!

  Compounded semiannually. P dollars is invested at annual interest rate r for 1 year. If the itnerest is compounded semiannually, then the polynomial p(1+r/2)^2 represents the value of the investment after 1 year. Rewrite this expression without paraenteses. Evaluatie the polynomial if P= $200 and r =10%.
: Hi! Pretty please with a cherry on the top would you help me solve this problem! Thank you so much. Appreciate it!

  Compounded semiannually. P dollars is invested at annual interest rate r for 1 year. If the itnerest is compounded semiannually, then the polynomial p(1+r/2)^2 represents the value of the investment after 1 year. Rewrite this expression without paraenteses. Evaluatie the polynomial if P= $200 and r =10%.

Answer by ankor@dixie-net.com(4684)

Compounded semi-annually. P dollars is invested at annual interest rate r for 1
year. If the interest is compounded semi-annually, then the polynomial p(1+r/2)^2
represents the value of the investment after 1 year.
:
Rewrite this expression without parenthesis.
A = p(1+r/2)^2
:
FOIL (1+(r/2))*(1+(r/2)) to get rid of the exponent
A = p *(1 + 2(r/2) + r^2/4)
:
A = p (1 + r + r^2/4); cancel 2
:
Multiply what's inside the brackets by p
A = p + pr + (pr^2)/4
:
Evaluate the polynomial if P= $200 and r =10%.
Substitute for p and r in the above equation
:
A = 200 + 200*.10 + (200*.1^2)/4
:
A = 200 + 20 + (200*.01)/4
:
A = 220 + 2/4
:
A = 220 + .50
:
A = 220.50
:
:
Check solution on a calc: enter 200*(1+(.1/2))^2 = 220.5